# Large Vision-Language Model Alignment and Misalignment: A Survey Through the Lens of Explainability

### v1

@@ 5.1 Standardized Benchmarks
5.2 Explainability based Diagnose
@@
5.3 Architectural Innovations
Current LVLM architectures face fundamental challenges, including significant ability gaps between visual encoders and LLMs, persistent attention imbalances between modalities, and knowledge conflicts between visual and linguistic representations. While most existing solutions focus on improving training procedures or adding post-processing steps, future research should focus on architectural innovations that address these structural limitations. This could include developing novel integration mechanisms that better balance the capabilities of visual and language components, and dynamic architectures that can adaptively adjust their attention mechanisms to maintain equilibrium between modalities. The field would benefit from multi-stage processing architectures that explicitly manage knowledge conflicts through specialized components for different levels of semantic understanding. Additionally, new transformer architectures specifically designed for vision-language tasks, rather than adapted from unimodal architectures, could help bridge the ability gap between visual and linguistic processing.
6 Conclusions
In this paper, we present a systematic survey of alignment and misalignment in LVLMs through an explainability lens. Our investigation demonstrates that achieving proper alignment involves complex interactions between data quality, model architecture, and inference procedures. We have developed a categorization of misalignment into object, attribute, and relational levels, offering a clear framework for understanding these challenges and developing targeted solutions. The examination of current mitigation strategies reveals a spectrum of approaches, from computationally intensive parameter-tuning methods to more practical parameter-frozen solutions, each with distinct trade-offs between effectiveness and implementation feasibility. Lastly, we have identified several key directions for future research, which will be essential for creating more reliable and capable vision-language systems that maintain robust alignment while serving diverse real-world applications.

### v2

@@ F.1 Standardized Benchmarks
F.2 Explainability based Diagnose
@@
F.3 Architectural Innovations
Current LVLM architectures face fundamental challenges, including significant ability gaps between visual encoders and LLMs, persistent attention imbalances between modalities, and knowledge conflicts between visual and textual representations. While most existing solutions focus on improving training procedures or adding post-processing steps, future research should focus on architectural innovations that address these structural limitations. This could include developing novel integration mechanisms that better balance the capabilities of visual and language components, and dynamic architectures that can adaptively adjust their attention mechanisms to maintain equilibrium between modalities. The field would benefit from multi-stage processing architectures that explicitly manage knowledge conflicts through specialized components for different levels of semantic understanding. Additionally, new transformer architectures specifically designed for vision-language tasks, rather than adapted from unimodal architectures, could help bridge the ability gap between visual and textual processing.
Figure 4: This figure presents examples of descriptions generated by different LVLMs when given the image shown in the upper left corner. The highlighted text segments in the descriptions represent different types of misalignment. Specifically, red highlights indicate object misalignment. Yellow highlights represent attribute misalignment. Blue highlights denote relational misalignment.
Figure 5: This figure presents examples of descriptions generated by different LVLMs when given the image shown in the upper left corner. The highlighted text segments in the descriptions represent different types of misalignment. Specifically, red highlights indicate object misalignment. Yellow highlights represent attribute misalignment. Blue highlights denote relational misalignment.
Figure 6: This figure presents examples of descriptions generated by different LVLMs when given the image shown in the upper left corner. The highlighted text segments in the descriptions represent different types of misalignment. Specifically, red highlights indicate object misalignment. Yellow highlights represent attribute misalignment. Blue highlights denote relational misalignment.